A 4.0 mg/dl standard solution is needed.

To prepare 100 ml of the working standard,

how much stock standard of a 1-mg/ml solution is needed?

I need this one using the specific information so I understand the V1C1 = V2C2

Please explain in detail how to answer the Question. Thank you
Answer #1

use dilution formula

C1*V1 = C2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

C1 = 1 mg/mL

C2 = 4.0 mg/dL = 4.0 mg/100 mL = 0.04 mg/mL

V2 = 100 mL

use:

C1*V1 = C2*V2

V1 = (C2 * V2) / C1

V1 = (0.04*100)/1

V1 = 4.0 mL

Answer: 4.0 mL